OpenStudy (deoxna):

f(x)=cos(x)+(1/2)cos(2x) Show that f(-x)=f(x) Is the fact that cos(-a)=cos(a) proof enough?

OpenStudy (thomas5267):

I think you still have to show that \(f(x)=f(-x)\).

OpenStudy (deoxna):

Ok so I write out: \[f(-x)=\cos(-x)+(1/2)\cos(-2x)\] then \[\cos(-a)=\cos(a)\] \[f(-x)=\cos(x)+(1/2)\cos(2x)\] I just can't see any other way of proving this, other than perhaps double angle identities, but then I would still have to rely on cos(x)=cos(-x) to prove the first term.

OpenStudy (thomas5267):

I think using \(\cos (x)=\cos (-a)\) will do.

OpenStudy (deoxna):

Ok, thank you. I wasn't sure about that one...

OpenStudy (thomas5267):

I think you can't use double angle identity to prove it though.

OpenStudy (deoxna):

Yeah, cause you'd still have (cos(2x)^2 which you can't really simplify...